OpenStudy (anonymous):

What is the area of a sector with a central angle of 160 degrees and a diameter of 5.8 m? Round the answer to the nearest tenth.

OpenStudy (anonymous):

first you gotta find the area of the circle

OpenStudy (mathstudent55):

Do you know how to find the area of a circle?

OpenStudy (anonymous):

A=pi(r^2)

OpenStudy (anonymous):

this area is for all 360. hene for unit angle at the centre the area must be pi(r^2) /360 thus for 160 as central angle

OpenStudy (anonymous):

you can have the sector are = 160 /360 *pi(r^2) now you know the radius so you can ...

OpenStudy (mathstudent55):

\(\LARGE A_{sector} = \dfrac{n}{360^o} \pi r^2 \) where n is the degree measure of the sector, and r is the radius of the circle.

OpenStudy (dengeki_daisy):

is the answer 11.7

OpenStudy (dengeki_daisy):

Qwertty123

OpenStudy (mathstudent55):

Yes, the area of the sector is 11.7 m^2.

OpenStudy (dengeki_daisy):

yeayyy ^-^

OpenStudy (mathstudent55):